OpenStudy (anonymous):

What is the solution to the following system of equations? y = x^2 + 5x + 3 y = x^2 + 3x – 3 A. (–3, 3) B. (5, 3) C. (5, –3) D. (–3, –3)

jhonyy9 (jhonyy9):

multiplie second equation by -1 and assum to first equations and will get y=x^2 +5x +3 -y=-x^2-3x +3 --------------- 0=0 + 2x +6 0=2x+6 2x=-6 x=-6/2 x=-3 y=x^2 +3x -3 y=(-3)^2 +3(-3) -3 y=9 -9 -3 y=-3 (-3,-3) so D. is the right answer